whats 500,000 times 500,000
Found 2 solutions by rwm, solver91311:
Answer by rwm(914)

5*10^5*5*10^5
=25*10^10=2.5*10^11

Answer by solver91311(24713)



5 times 5 is 25, then there are 5 zeros in one of your factors and 5 zeros in the other factor, so a total of 10 zeros. Hence, 250 Billion.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism